# Solving the Optimal Experiment Design Problem with Mixed-Integer Convex Methods

### v5

@@ 0.A.1 Lipschitz smoothness of the Fusion Problems
Let us now prove $L$-smoothness for the Fusion Problems under the A-criterion and D-criterion.
Theorem 0.A.1
The functions $f_{F}(\mathbf{x}):=-\log\det\left(X_{C}(\mathbf{x})\right)$ and $g_{F}(\mathbf{x}):=\operatorname{Tr}\left(\left(X_{C}(\mathbf{x})\right)^{-p}\right)$ for $p\in{\mathbb{R}}_{>0}$ are $L$-smooth with Lipschitz constants
$L_{f}=\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\|A\|_{2}^{2}}{\lambda_{\min}(C)^{2}}$
and
$L_{g}=p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(C)^{2+p}},$
respectively.
Proof
For better readability and simplicity, we ignore the dependency on $\mathbf{x}$ for now. As stated above, we can prove $L$-smoothness by showing that the largest eigenvalue of the Hessians is bounded for any $\mathbf{x}$ in the domain. We have
$\nabla^{2}f(\mathbf{x})$ $=(AA^{\intercal})\circ(AX_{C}^{-2}A^{\intercal})$
and
$\nabla^{2}g(\mathbf{x})$ $=p(p+1)\left(AA^{\intercal}\right)\circ\left(AX_{C}^{-p-2}A^{\intercal}\right).$
Observe that the Hessian expressions are very similar. Hence, we will only present the proof for $g$. The proof of $f$ follows the same argumentation. By (Horn, 1990, page 95), it is known that the maximum eigenvalue of a Hadamard product between two positive semi-definite matrices $N$ and $M$, $\lambda_{\max}(N\circ M)$, is upper-bounded by the maximum diagonal entry of $N$ and the largest eigenvalue of $M$, so
$\lambda_{\max}(N\circ M)$ $\leq\max_{i}N_{ii}\,\lambda_{\max}(M).$ (3)
In our case, $N=AA^{\intercal}$ and $M=AX_{C}^{-p-2}A^{\intercal}$. By Corollary 1 we have that $X_{C}^{-p-2}$ is positive definite. Thus, both factors are semi-definite and thus we may use the provided bound:
$\lambda_{\max}\left(\nabla^{2}g(\mathbf{x})\right)$ $\leq\max\limits_{i}(AA^{\intercal})_{ii}\,\lambda_{\max}(AX_{C}^{-p-2}A^{\intercal}).$
Note that the first factor is simply:
For the second factor, we can exploit positive semi-definiteness with the spectral norm:
$\leq\left\|A\right\|_{2}^{2}\left\|X_{C}^{-p-2}\right\|_{2}.$
By Lemma 1, we have:
$\lambda(A^{p})$ $=\lambda(A)^{p}$
$=\left\|A\right\|_{2}^{2}\lambda_{\max}(X_{C}^{-1})^{2+p}$
$=\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(X_{C})^{2+p}}.$
By the definition of $X$ and by Weyl’s inequality on the eigenvalues of sums of positive semi-definite matrices (Weyl, 1912), we have that:
Hence,
Combining (3) with (4) and (5) yields
$\lambda_{\max}\left(\nabla^{2}g(\mathbf{x})\right)$ $\leq p(p+1)\left(\max\limits_{i}v_{i}^{\intercal}v_{i}\right)\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(C)^{2+p}}.$ (6)
This concludes the proof. ∎
Note that the proof works because we can bound the minimum eigenvalue of $X_{C}$ away from zero thanks to the positive definite matrix $C$. This is not the case for the information matrices $X(\mathbf{x})$ of the Optimal Problems as they can be singular.
0.A.2 Generalized Self-Concordance
Lacking $L$-smoothness on the feasible region for the Optimal Problems, we need a different property guaranteeing convergence. As stated in Section 3, we need to prove that the objective function resulting from the A-criterion is generalized self-concordant to establish convergence of Frank-Wolfe. In fact, we can prove the statement for the General-Trace-Inverse (GTI) $g(x)=\operatorname{Tr}\left(X^{-p}\right)$ for $p>0$.
Let $A\in\mathbb{S}_{++}^{n},B\in\mathbb{S}_{+}^{n}$ and $t\in{\mathbb{R}}$ such that $A+tB\in\mathbb{S}_{++}^{n}$. We define $h(t)=g(A+tB)$. For the proof of self-concordance, we need the first three derivatives of $h(t)$.
$h(t)$ $=\operatorname{Tr}((A+tB)^{-p})=\operatorname{Tr}\left(\left((A+tB)^{-1}\right)^{p}\right)$
Using Lemma 3 for the first derivative, yields:
@@ 0.A.2 Generalized Self-Concordance
$=-p\operatorname{Tr}\left(\left((A+tB)^{-1}\right)^{p}B(A+tB)^{-1}\right).$
For the second derivative, we find
@@
$=p\operatorname{Tr}\left((p+1)\left((A+tB)^{-1}\right)^{p}B(A+tB)^{-1}B(A+tB)^{-1}\right).$
Restricting $h^{\prime\prime}$ to $t=0$, yields:
@@
Lastly, for the third derivative, we have:
@@
And restricted to $t=0$, yields:
$h^{\prime\prime\prime}(t)_{\restriction t=0}$ $=-p(p+1)(p+2)\operatorname{Tr}\left(A^{-p}BA^{-1}BA^{-1}BA^{-1}\right).$
Proof (Proof of Theorem 3.1)
We define $M(t)=V+tU$ where $V\in\mathrm{dom}(g)$, $U\in\mathbb{S}_{+}^{n}$ and $t\in{\mathbb{R}}$ such that $V+tU\in\mathbb{S}_{++}^{n}$. Then:
$h(t)$ $=g(M(t))=\operatorname{Tr}\left(M(t)^{-p}\right).$
From the derivative computation above, we have
Thus, the condition we want to satisfy is:
$p(p+1)(p+2)\operatorname{Tr}\left(V^{-p}UV^{-1}UV^{-1}UV^{-1}\right)$ $\leq M_{f}p(p+1)Tr\left(V^{-p}UV^{-1}UV^{-1}\right)^{\nu/2}.$
Note that
$\operatorname{Tr}\left(V^{-p}UV^{-1}UV^{-1}UV^{-1}\right)=\operatorname{Tr}\left(V^{-\frac{p+1}{2}}UV^{-1}UV^{-1}UV^{-\frac{p+1}{2}}\right)\geq 0$
and
$\operatorname{Tr}\left(V^{-p}UV^{-1}UV^{-1}\right)=\operatorname{Tr}\left(V^{-\frac{p+1}{2}}UV^{-1}UV^{-\frac{p+1}{2}}\right)\geq 0$
by positive definiteness of $V$ and semi-definiteness $U$. If the trace is equal to 0, the statement holds directly. Otherwise, we divide both sides by the LHS and first try to lower bound the following fraction:
$\frac{\operatorname{Tr}\left(V^{-\frac{p+1}{2}}UV^{-1}UV^{-\frac{p+1}{2}}\right)^{\nu/2}}{\operatorname{Tr}\left(V^{-\frac{p+1}{2}}UV^{-1}UV^{-1}UV^{-\frac{p+1}{2}}\right)}.$
We want to lower-bound this fraction. To that end, we define $H=V^{-1/2}UV^{-1/2}$ and $L=V^{-p/2}H$, then:

### v6

@@ A.1 General power of a matrix
Lemma A.21.
For $X\in\mathbb{S}_{++}^{n}$, we have $X^{-p}=\left(X^{-1}\right)^{p}$.
Proof.
$\left(X^{-1}\right)^{p}$ $=\sum\limits_{k=0}^{\infty}\frac{p^{k}}{k!}\log\left(X^{-1}\right)^{k}$
$\log\left(X^{-1}\right)$ $=V\log(D^{-1})V^{\intercal}=-V\log(D)V^{\intercal}=-\log(X)$
$\left(X^{-1}\right)^{p}$ $=\sum\limits_{k=0}^{\infty}\frac{p^{k}}{k!}\left(-\log(X)\right)^{k}$
$=\sum\limits_{k=0}^{\infty}\frac{(-p)^{k}}{k!}\log(X)^{k}$
$=X^{-p}$
Remark.
On the contrary, $\left(X^{p}\right)^{-1}\neq X^{-p}$ in general.
A.2 Convexity of the objectives
In this subsection, we present the convexity proofs for all the considered objectives. Note that concavity of the $\log\det$ is well known and is stated here for completeness.
Lemma A.22.
The function $f(X)=-\log\det(X)$ is convex for $X\in\mathbb{S}_{++}^{n}$.
Proof.
Let $A\in\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}$ and $t\in{\mathbb{R}}$ such that $A+tB\in\mathbb{S}_{++}^{n}$. We define $h(t)=-\log\det\left(A+tB\right)$.
$-\log\det(A+tB)$ $=-\log\det\left(A^{\frac{1}{2}}A^{\frac{1}{2}}+tA^{\frac{1}{2}}A^{-\frac{1}{2}}BA^{-\frac{1}{2}}A^{\frac{1}{2}}\right)$
$=-\log\det\left(A^{\frac{1}{2}}\left(I+tA^{-\frac{1}{2}}BA^{-\frac{1}{2}}\right)A^{\frac{1}{2}}\right)$
$=-\log\det(A)-\log\det\left(I+tA^{-\frac{1}{2}}BA^{-\frac{1}{2}}\right)$
Observe that $I+tA^{-\frac{1}{2}}BA^{-\frac{1}{2}}$ is a positive definite matrix and its eigenvalues are $1+t\lambda_{i}$ with $\lambda_{i}$ being the $i$-th eigenvalue of $A^{-\frac{1}{2}}BA^{-\frac{1}{2}}$. Note that this also means $01+t\lambda_{i}>$ 0.
$=-\log\det(A)-\log\left(\prod\limits_{i=1}^{n}(1+t\lambda_{i})\right)$
$=-\log\det(A)-\sum\limits_{i=1}^{n}\log\left(1+t\lambda_{i}\right)$
The first and second derivatives of $g$ are
$h^{\prime}(t)$ $=-\sum\limits_{i=1}^{n}\frac{1}{1+t\lambda_{i}}\lambda_{i}$
$h^{\prime\prime}(t)$ $=\sum\limits_{i=1}^{n}\frac{\lambda_{i}^{2}}{\left(1+t\lambda_{i}\right)}$
$>0\quad\forall t$
Thus, $f(X)$ is convex. ∎
Lemma A.23.
The function $g(X)=\operatorname{Tr}(X^{-p})$ is convex for $X\in\mathbb{S}_{++}^{n}$ and $p>0$.
Proof.
Let $A\in\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}$ and $t\in{\mathbb{R}}$ such that $A+tB\in\mathbb{S}_{++}^{n}$. We define $h(t)=\operatorname{Tr}\left(\left(A+tB\right)^{-p}\right)$.
$h(t)$ $=\operatorname{Tr}((A+tB)^{-p})$
$=\operatorname{Tr}\left(\left((A+tB)^{-1}\right)^{p}\right)$
Using Lemma A.20 for the derivative yields
@@ A.2 Convexity of the objectives
$=-p\operatorname{Tr}\left(\left((A+tB)^{-1}\right)^{p}B(A+tB)^{-1}\right)$
@@
$=p\operatorname{Tr}\left((p+1)\left((A+tB)^{-1}\right)^{p}B(A+tB)^{-1}B(A+tB)^{-1}\right)$
Restricting $h^{\prime\prime}$ to $t=0$ yields
@@
For the third derivative, we have
$h^{\prime\prime\prime}(t)$ $=p(p+1)\operatorname{Tr}\left(p\left((A+tB)^{-1}\right)^{p-1}(-1)(A+tB)^{-1}B(A+tB)^{-1}B(A+tB)^{-1}B(A+tB)^{-1}\right.$
[…]
@@
$h^{\prime\prime\prime}(t)_{\restriction t=0}$ $=-p(p+1)(p+2)\operatorname{Tr}\left(A^{-p}BA^{-1}BA^{-1}BA^{-1}\right)$
By Corollary A.19, $A^{-p}$ is positive definite and thus, has a unique root $A^{-p/2}$. Using the traces cyclic property, we find
$h^{\prime\prime}(t)_{\restriction t=0}$ $=p(p+1)\operatorname{Tr}\left(A^{-\frac{p+1}{2}}BA^{-1}BA^{-\frac{p+1}{2}}\right)$
$=p(p+1)\sum\limits_{i=1}^{n}\lambda_{i}\left(A^{-\frac{p+1}{2}}BA^{-1}BA^{-\frac{p+1}{2}}\right)$
$\geq 0$
The last inequality holds because the matrix $A^{-\frac{p+1}{2}}BA^{-1}BA^{-\frac{p+1}{2}}$ is positive definite by positive definiteness of $A$, $B$ and $A^{-p}$. Thus, $g(X)$ is convex. ∎
Lemma A.24.
The function $k(X)=\log\left(\operatorname{Tr}(X^{-p})\right)$ is convex on $\mathbb{S}_{++}^{n}$ for $p>0$.
Proof.
Let $A\in\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ s.t. $A+tB\in\mathbb{S}_{++}^{n}$. We define $h(t)=k(A+tB)=\log\left(\operatorname{Tr}((A+tB)^{-p})\right)$.
$h(t)$ $=\log\left(\operatorname{Tr}\left((A+tB)^{-p}\right)\right)$
$h^{\prime}(t)$ $=-\frac{p}{\operatorname{Tr}((A+tB)^{-p})}\operatorname{Tr}((A+tB)^{-p-1}B)$
$h^{\prime\prime}(t)$ $=p\frac{(p+1)\operatorname{Tr}\left((A+tB)^{-p-2}BB\right)\operatorname{Tr}\left((A+tB)^{-p}\right)-p\operatorname{Tr}\left((A+tB)^{-p-1}B\right)^{2}}{\operatorname{Tr}\left((A+tB)^{-p}\right)^{2}}$
Restricting us to zero, we get
$h^{\prime\prime}(t)_{\restriction t=0}$ $=p\frac{(p+1)\operatorname{Tr}\left(A^{-p-2}BB\right)\operatorname{Tr}\left(A^{-p}\right)-p\operatorname{Tr}\left(A^{-p-1}B\right)^{2}}{\operatorname{Tr}\left(A^{-p}\right)^{2}}.$